a mixture of CH4 and C2H2 occupied a certain volume at a total pressure of 63mm.the sample was burnt to CO2 and H2O and the CO2 alone was collected and its pressure was found to be 69mm in the same volume and at the same temperature as the original mixture .what fraction of the mixture was methane.

Grade:upto college level

2 Answers

Sunil Kumar FP
9 years ago
CH4 +2O2 ----CO2 +2H20
C2H2 +5/2O2 ----2CO2 +H20

let the partial pressure of CH4 be p thus the partial pressure of C2H2 is 63-p
thus following the question
total pressure of CO2 after combustion=69 mm
therefore
p +2(63-p) =69
p=57 mm
partial pressure of C2H2=6mm

percentage of CH4=57/63*100 =90.41%
Jk
3 years ago
Ch4:0.90
As we have to find fraction 
P+2(p-63)=69
We can get 
P is 57 
Then ans
As long as you can do it for me to come over and over again and I don't know what I want to say that I have to be a good Day World Series of poker face and I don't know what I want to say that I have to